# Deep Reinforcement Learning for Dynamic Order Picking in Warehouse Operations

### v1

@@ Abstract
Order picking is a crucial operation in warehouses that significantly impacts overall efficiency and profitability. This study addresses the dynamic order picking problem, a significant concern in modern warehouse management where real-time adaptation to fluctuating order arrivals and efficient picker routing are crucial. Traditional methods, often assuming fixed order sets, fall short in this dynamic environment. We utilize Deep Reinforcement Learning (DRL) as a solution methodology to handle the inherent uncertainties in customer demands. We focus on a single-block warehouse with an autonomous picking device, eliminating human behavioral factors. Our DRL framework enables the dynamic optimization of picker routes, significantly reducing order throughput times, especially under high order arrival rates. Experiments demonstrate a substantial decrease in order throughput time and unfulfilled orders compared to benchmark algorithms. We further investigate integrating a hyperparameter in the reward function that allows for flexible balancing between distance traveled and order completion time. Finally, we demonstrate the robustness of our DRL model for out-of-sample test instances.
@@ 1 Introduction
Order picking, the process of retrieving specific items from warehouse storage to fulfill customer orders, is a labor-intensive and time-consuming operation. Order pickers traverse the warehouse, collecting items from various locations before delivering them to a designated depot. Notably, order picking can constitute about 50% of the total order processing time (Charkhgard and Savelsbergh, 2015; Silva et al., 2020). Recent growth in e-commerce, coupled with heightened customer expectations for fast delivery and increased market competition, has escalated the demand for rapid order fulfillment (Marchet et al., 2015; Van Gils et al., 2018; Rasmi et al., 2022). Efficient order picking, which necessitates solving a routing problem, is crucial for several reasons. It directly enhances operational efficiency, leading to increased order throughput within the warehouse and improved customer satisfaction (Giannikas et al., 2017). Additionally, streamlining the picking process translates to cost reductions for the business.
@@
In modern warehouse operations, the ability to accurately forecast demand and dynamically adjust order picking decisions is paramount (Lu et al., 2016; Dauod and Won, 2022). However, the existing literature predominantly focuses on exact and heuristic methods that assume a fixed set of orders is known in advance. While this approach is suitable for static scenarios, it often fails to effectively address the dynamic nature of real-world warehouse operations. To overcome these challenges, a dynamic and adaptive approach to demand forecasting and picker routing is essential. This would enable warehouses to respond swiftly to real-time order arrivals, optimizing routing decisions, minimizing operational costs, and maximizing overall efficiency.
Our research investigates picker-to-parts systems within single-block warehouses, focusing on how an autonomous picker navigates the aisles to retrieve items. This focus aligns with the growing trend of warehouse automation (Roy et al., 2019; Winkelhaus et al., 2021), and we assume the picking device flawlessly execute recommended routing decisions, thereby eliminating behavioral factors associated with human pickers. We demonstrate that a dynamic, single-picker order picking method can substantially reduce picking time in this setting. Our study serves as a foundation for future research to explore the complexities of dynamic order picking in multi-block warehouses with multiple coordinated pickers.
@@
We employ a Deep Reinforcement Learning (DRL) framework for addressing the dynamic order picking problem due to its inherent ability to handle sequential decision-making under uncertainty. This approach is particularly well-suited for the real-world warehouse environment (Begnardi et al., 2024), characterized by the constantly changing nature of order arrivals, item locations, picker location, and picker availability. By leveraging DRL’s capability to learn and adapt in real-time, and forecast order arrivals to optimize picker routes dynamically, we can significantly improve warehouse efficiency and throughput. We demonstrate that traditional deep neural network architectures are sufficient to achieve this goal, as opposed to more complex state-of-the-art DRL models, which require longer training times or heavier computational resources.
@@
Our experiments demonstrate that as order arrival rates increase, relying solely on the shortest-distance for order picking becomes inefficient. In scenarios with high order frequencies, the policies learned by our DRL agent substantially enhances the order fulfillment efficiency. For example, when orders follow a Poisson process with an arrival rate ($\lambda$) of 0.09, our approach reduces order throughput time by a substantial 420% compared to existing benchmark algorithms. Our approach maintains an unfulfilled order rate of approximately 2% within the work shift, while the benchmark algorithms result in up to 18% of orders remaining unfulfilled. Additionally, we demonstrate that incorporating a hyperparameter ($\alpha$) into the reward function allows us to effectively balance the trade-off between distance traveled and order throughput times, aligning with the preferences of decision-makers.
@@ 2 Literature Review
In this section, we first review the traditional optimization approaches for order picker routing in warehouses. We then discuss the application of DRL in addressing various warehouse operational challenges, demonstrating its potential for intelligent decision-making in complex environments. Finally, we identify a critical research gap within the dynamic order picking literature, highlighting the need for our proposed DRL-based approach.
2.1 Optimization Approaches for Order Picking Routing
Numerous methods have been proposed in the literature for order picker routing. However, most methods focus on static problem scenarios. A notable contribution for single-picker routing in single-block warehouses is Ratliff and Rosenthal (1983), who developed an efficient algorithm leveraging the graph-based structure of order storage locations in the warehouse. This algorithm has been extended to accommodate different warehouse layouts, including two-block (Roodbergen and De Koster, 2001), fish-bone (Çelk and Süral, 2014), and chevron (Masae et al., 2020b) configurations. The mathematical formulations for the same problem are presented in Scholz et al. (2016) and Pansart et al. (2018), offering advantages over the standard Traveling Salesperson Problem (TSP) solvers by utilizing the warehouse structure properties to eliminate redundant possibilities inherent in the TSP. However, these approaches are primarily designed for static environments and lack the flexibility to accommodate dynamic order arrivals, a common occurrence in real-world operations. Furthermore, practical implementation often necessitates an order batching policy due to picker capacity constraints. Lu et al. (2016) addressed these limitations by employing a first-come-first-served order batching policy and extending the static algorithm to dynamic scenarios. By utilizing this interventionist routing approach, the authors demonstrate a significant reduction in average order completion times compared to both heuristics and the static exact algorithm.
Heuristic algorithms are a popular focus in warehouse operations research due to their simplicity and ease of implementation. For example, the S-shape heuristic, where pickers traverse aisles in an S-shape curve, offers a straightforward routing policy. Hall (1993) provides a comparative analysis of various heuristic routing methods based on the number of pick locations within an aisle. The largest gap heuristic generally outperforms others when the pick list is small. However, the inherent nature of heuristic algorithms often leads to suboptimal decisions, potentially impacting overall system efficiency. Lu et al. (2016) demonstrated that their dynamic programming algorithm outperforms the largest gap heuristic in terms of both order completion time and total travel distance.
Order batching is a critical strategy that can significantly reduce the distance traveled by order pickers (De Koster et al., 2007; Henn, 2012; Muter and Öncan, 2022). However, most studies focus on either order batching or routing strategies, often employing simplistic approaches for the other aspect. Scholz and Wäscher (2017) introduced an iterated search heuristic to address the combined problem of order batching and picker routing. Their approach demonstrated better performance compared to other common methods, highlighting the advantages of integrating these two warehouse optimization problems. Recent research has seen a growing interest in dynamic pick list updating, notably in Dauod and Won (2022) and Yang et al. (2021). Due to limitations in scope, we benchmark our proposed method against Lu et al. (2016), rather than conducting a comprehensive comparison with other heuristics. Importantly, our proposed approach is dynamic and produces decisions almost instantaneously during testing, while the neighborhood searches involved in heuristics are time-consuming, leading to cumulative wait times in a dynamic order picking scenario.
For a more comprehensive literature review on order picker routing, including a detailed overview of various operational procedures, readers may refer to De Koster et al. (2007), Van Gils et al. (2018), and Masae et al. (2020a).
@@ 2.2 Deep Reinforcement Learning for Warehouse Operations
Deep Reinforcement Learning (DRL) has garnered increasing attention in the warehousing domain, with recent studies primarily focusing on order batching and assignment problems. Cals et al. (2021) pioneered DRL in warehousing by addressing the order batching and sequencing problem, aiming to minimize tardy orders. Later, Beeks et al. (2022) extended this method to a multi-objective domain, considering picking costs and analyzing the trade-off between order tardiness and picker efficiency. Li et al. (2019) proposed a DRL architecture for assigning orders to autonomous robots in a warehouse to reduce makespan in the event of traffic conflicts. Notably, these methods do not directly address optimal picker routing with the objective of minimizing travel distance or order completion times. Additionally, some studies utilize DRL for warehouse inventory management (Kaynov et al., 2024; Tian et al., 2024).
@@
Dunn et al. (2024) address the picker routing problem using a DRL approach for combinatorial optimization. Their model, utilizing an attention-based neural network, generates high-quality solutions and is particularly effective when extended to large-scale problems where exact methods may be computationally expensive. The authors demonstrate the effectiveness of their DRL model compared to the benchmark algorithm of Ratliff and Rosenthal (1983). However, this study is confined to the static problem variant. In a recent study, Begnardi et al. (2024) propose multiple graph neural network-based DRL models to solve the collaborative human-robot picking problem, formulated as an online bipartite matching problem. However, their evaluation is limited to small-scale grid-based instances with several pickers, and benchmarks are solely compared against variations of the greedy heuristic. Given the large number of pickers within their small-scale test instances, the study focuses more on the order assignment problem rather than the routing problem.
@@ 2.3 Research Gap
Our study is the first to comprehensively address the dynamic order picker routing problem using DRL, where the agent implicitly learns order batching strategies. This approach integrates order assignment and routing within the RL agent’s learning environment. The RL agent’s ability for rapid decision-making during testing is vital to prevent unnecessary delays in picker routing. We prioritize the development of a simple yet effective model for single-block warehouses and demonstrate its superior performance in minimizing order processing delays.
@@ 3 Problem Description
Consider the order-picking problem for a single picker within a rectangular warehouse. The warehouse consists of a single block with multiple parallel aisles, interconnected by two cross-aisles at the ends. Inventory storage locations are situated on either side of each aisle. Figure 1 presents a visual representation of this warehouse layout. The picker must collect the requested items from designated storage locations within the warehouse. The picker can move through the aisles to access storage locations on both sides, and we assume that there is no horizontal travel distance within an aisle. To traverse between different aisles, the picker must utilize one of the two cross-aisles. These cross-aisles do not contain any storage locations; they solely serve as pathways for movement between the aisles. Note that the picker can only enter or exist an aisle through one of the two cross-aisles.
@@
We assume that the requested items are identical in size and hold the same level of priority. The order-picking device or vehicle utilized for transporting these items has a total capacity of $K$ items. Initially, the picker starts from the depot and moves through the warehouse to collect the assigned items and delivers them to the depot, subject to the capacity constraint at all times. Figure 1 illustrates that this warehouse layout offers two alternative paths the picker can take to collect the next assigned order, denoted by the distances $\rho$. We will discuss the notation details of distances in the next section. The primary objective is to minimize the average waiting time for orders, and thereby increasing the throughput in warehouse operations. The picker route is determined based on the demand forecast, and a dynamically updated pick-list, which is adjusted based on incoming orders and the remaining orders to be collected. This optimization problem is considered over a time horizon that resembles a human picker’s work shift.
In most warehouse order-picking scenarios, customer demand fluctuates throughout a work shift, unlike static problems with pre-defined pick lists. This dynamic environment requires real-time decision-making regarding order batching and traversal routes to minimize order wait times. Reducing these wait times not only enhances customer satisfaction but also frees up resources for fulfilling additional orders, especially critical in the growing field of e-commerce. Our approach further improves order picking by incorporating implicit demand forecasting, allowing pickers to anticipate upcoming orders and optimize their paths. This contrasts with traditional methods based solely on observed order arrivals. Our findings demonstrate that this dynamic, priority-driven behavior leads to significant reductions in order wait times.

### v2

@@ Abstract
Order picking is a pivotal operation in warehouses that directly impacts overall efficiency and profitability. This study addresses the dynamic order picking problem, a significant concern in modern warehouse management, where real-time adaptation to fluctuating order arrivals and efficient picker routing are crucial. Traditional methods, which often depend on static optimization algorithms designed around fixed order sets for the picker routing, fall short in addressing the challenges of this dynamic environment. To overcome these challenges, we propose a Deep Reinforcement Learning (DRL) framework tailored for single-block warehouses equipped with an autonomous picking device. By dynamically optimizing picker routes, our approach significantly reduces order throughput times and unfulfilled orders, particularly under high order arrival rates. We benchmark our DRL model against established algorithms, utilizing instances generated based on standard practices in the order picking literature. Experimental results demonstrate the superiority of our DRL model over benchmark algorithms. For example, at a high order arrival rate of 0.09 (i.e., 9 orders per 100 units of time on average), our approach achieves an order fulfillment rate of approximately 98%, compared to the 82% fulfillment rate observed with benchmarking algorithms. We further investigate the integration of a hyperparameter in the reward function that allows for flexible balancing between distance traveled and order completion time. Finally, we demonstrate the robustness of our DRL model on out-of-sample test instances.
@@ 1 Introduction
Order picking, the process of retrieving specific items from warehouse storage to fulfill customer orders, is a labor-intensive and time-consuming operation. Order pickers traverse the warehouse, collecting items from various locations before delivering them to a designated depot. Notably, order picking can constitute about 50% of the total order processing time (Charkhgard and Savelsbergh, 2015; Silva et al., 2020). Recent growth in e-commerce, coupled with heightened customer expectations for fast delivery and increased market competition, has escalated the demand for rapid order fulfillment (Marchet et al., 2015; Rasmi et al., 2022; Van Gils et al., 2018). Efficient order picking, which necessitates solving a routing problem, is crucial for several reasons. It directly enhances operational efficiency, leading to increased order throughput within the warehouse and improved customer satisfaction (Giannikas et al., 2017). Additionally, streamlining the picking process translates to cost reductions for the business.
@@
In modern warehouse operations, the ability to accurately forecast demand and dynamically adjust order picking decisions is paramount (Dauod and Won, 2022; Lu et al., 2016). However, the existing literature predominantly focuses on exact and heuristic methods that assume a fixed set of orders is known in advance. While this approach is suitable for static scenarios, it often fails to effectively address the dynamic nature of real-world warehouse operations. To overcome these challenges, a dynamic and adaptive approach to demand forecasting and picker routing is essential. This would enable warehouses to respond swiftly to real-time order arrivals, optimizing routing decisions, minimizing operational costs, and maximizing overall efficiency.
@@
This research investigates picker-to-parts systems within single-block warehouses, focusing on how an autonomous picker navigates the aisles to retrieve items. This focus aligns with the growing trend of warehouse automation (Bansal and Roy, 2021; Mirzaei et al., 2021; Roy et al., 2019; Winkelhaus et al., 2021, 2022) and assumes that the picking device flawlessly executes the recommended routing decisions, thereby eliminating behavioral factors associated with human pickers. To address the order picking problem, we employ a Deep Reinforcement Learning (DRL) framework, well-suited for its ability to handle sequential decision-making under uncertainty. This approach is particularly well-suited for the real-world warehouse environment (Begnardi et al., 2024; Charkhgard et al., 2025), characterized by the constantly changing nature of order arrivals, item locations, picker location, and picker availability. By leveraging DRL’s ability to learn and adapt to the environment while dynamically forecasting order arrivals and optimizing picker routes, we can achieve significant improvements in warehouse efficiency and throughput. Additionally, we show that traditional deep neural network architectures are sufficient to meet these goals, avoiding the complexity, longer training times, and higher computational demands of state-of-the-art DRL models.
Our main contribution is a DRL framework that integrates order assignment and picker routing within a unified learning environment. By incorporating implicit batching strategies and demand forecasting, this approach effectively optimizes dynamic picker routing in single-block warehouses. We demonstrate that our dynamic order-picking method significantly reduces order completion times and provide benchmarking against established algorithms from the literature. Our study serves as a foundation for future research to explore the complexities of real-time order picking in multi-block warehouses with multiple coordinated pickers.
@@
Our experiments demonstrate that as order arrival rates increase, relying solely on the shortest-distance for order picking becomes inefficient. In scenarios with high order frequencies, the policies learned by our DRL agent substantially enhances the order fulfillment efficiency. For example, when orders follow a Poisson process with an arrival rate ($\lambda$) of 0.09, our approach reduces order throughput time by a substantial 420% compared to existing benchmark algorithms. Our approach maintains an unfulfilled order rate of approximately 2% within the work shift, while the benchmark algorithms result in up to 18% of orders remaining unfulfilled. We also incorporate a hyperparameter ($\alpha$) into the reward function, enabling effective balancing of the trade-off between distance traveled and order throughput times, thus aligning with the preferences of decision-makers.
@@ 2 Literature Review
In this section, we review the literature on traditional optimization methods and DRL applications for order batching and picker routing in warehouses, emphasizing their influence on our research design and methodological choices. Traditional optimization techniques, including exact and heuristic methods, form the foundation upon which our problem definition and benchmarks are based. Recent DRL-based approaches in warehouse operations highlight the potential for learning-based solutions in dynamic environments, directly motivating our adoption of DRL as a solution method. Additionally, we briefly explore DRL applications in related logistics domains characterized by dynamic task allocation and routing, reinforcing the suitability of DRL for addressing dynamic order picking challenges. Finally, we identify a significant research gap regarding real-time integration of dynamic batching and picker routing, situating our proposed DRL-based framework as a meaningful advancement in the field.
2.1 Optimization Approaches for Order Picking
Numerous optimization techniques have been developed in the literature for order picking, but these generally address static problem settings. A notable contribution is by Ratliff and Rosenthal (1983), who designed an efficient algorithm that leverages the warehouse graph structure to optimize single-picker routing in single-block layouts. Later adaptations of this approach accommodate diverse warehouse configurations, such as two-block (Roodbergen and De Koster, 2001), fish-bone (Çelk and Süral, 2014), and chevron layouts (Masae et al., 2020b). Alternatively, Scholz et al. (2016) and Pansart et al. (2018) offer mathematical models that outperform standard Traveling Salesperson Problem (TSP) solvers by eliminating redundant routes, leveraging the warehouse structure to optimize routing. However, these static methods lack adaptability to dynamic order arrivals, a key characteristic of real-world applications.
Addressing the challenges of dynamic warehouse environments, Lu et al. (2016) introduced a first-come-first-served order batching policy that adapts the static algorithm to handle dynamic order inflows. Their experiments indicate notable reductions in average order completion times compared to both static exact algorithm and traditional heuristics. This adaptation is a precedent for dynamic routing solutions, and forms a baseline for our proposed DRL method to further improve efficiency in dynamic warehouse settings.
Heuristic algorithms have received significant attention due to their simplicity and ease of implementation. The S-shape heuristic, for example, offers a straightforward routing policy by assigning picker routes through aisles in an S-shaped curve. Hall (1993) presents a comparative analysis of heuristic methods, showing the effectiveness of the largest-gap heuristic, particularly with smaller pick lists. Meta-heuristic algorithms have also been explored in the literature, though they often require careful adaptation and specific local search strategies to suit different warehouse designs. For instance, Schrotenboer et al. (2017) presents a hybrid genetic algorithm to determine routes that simultaneously handle the pickup of products in response to customer orders and delivery of returned products to storage locations. However, the nature of heuristic algorithms can lead to suboptimal outcomes, which may reduce overall system efficiency (Roodbergen, 2001). Notably, Lu et al. (2016) demonstrated that their dynamic programming algorithm outperforms the largest-gap heuristic in terms of both order completion time and total travel distance, justifying our choice to benchmark our model against Lu et al. (2016).
Order batching is another critical strategy that can significantly reduce the distance traveled by order pickers (De Koster et al., 2007; Henn, 2012; Muter and Öncan, 2022). Menéndez et al. (2017) employ a Variable Neighborhood Search (VNS) to group warehouse orders into picker-managed batches. However, most studies treat batching and routing as separate problems, simplifying one to optimize the other. An exception is Scholz and Wäscher (2017), who propose an integrated approach that combines batching with picker routing, demonstrating superior results. In Aerts et al. (2021), order batching and routing are tackled as a static vehicle routing problem, and designs a VNS heuristic for the warehouse environment. Recent research has seen a growing interest in dynamic pick list updating, notably in Dauod and Won (2022) and Yang et al. (2021). Despite these advances, heuristics for dynamic scenarios still pose challenges, as the iterative nature of neighborhood searches can lead to computational delays. Our method leverages DRL for real-time decision-making, effectively eliminating cumulative wait times often caused by the computational time associated with optimization approaches. For a more comprehensive literature review on order picker routing and warehouse operational procedures, readers may refer to De Koster et al. (2007), Van Gils et al. (2018), and Masae et al. (2020a).
@@ 2.2 Deep Reinforcement Learning for Warehouse Operations
Deep Reinforcement Learning (DRL) has garnered increasing attention in the warehousing domain, with recent studies primarily focusing on order batching and assignment problems. Cals et al. (2021) pioneered this approach, using DRL to minimize tardy orders in order batching. Subsequently, Beeks et al. (2022) extended this method to a multi-objective problem, considering picking costs and analyzing the trade-off between order tardiness and picker efficiency. Similarly, Li et al. (2019) employed DRL for assigning tasks to autonomous robots, aiming to minimize makespan in the event of traffic conflicts. Some recent studies also utilize DRL for warehouse inventory management (Kaynov et al., 2024; Tian et al., 2024). These studies demonstrate the potential of DRL, yet they primarily address order batching rather than the intricacies of dynamic picker routing by minimizing travel distances or order completion times.
@@
In picker routing, Dunn et al. (2024) introduced a DRL model for the order picking as a combinatorial optimization problem. The authors employ an attention-based neural network that generates high-quality solutions and proves particularly effective for large-scale problems where exact methods become computationally expensive. However, this approach is limited to static scenarios. Echoing the need for real-world applications, Roodbergen and Vis (2009) advocates dynamic models, underscoring the relevance of our research work. More recently, Begnardi et al. (2024) applied graph neural networks in DRL for collaborative human-robot picking, formulated as an online bipartite matching problem. However, their evaluation is limited to grid-based warehouse layouts with several pickers, and benchmarks are solely variations of the greedy heuristic. Given the large number of pickers within their small-scale test instances, the study places greater emphasis on the order assignment problem rather than the routing problem.
We would like to acknowledge other recent studies in the literature, such as Cheng et al. (2024) and Neves-Moreira and Amorim (2024), which illustrate the flexibility of DRL in targeted warehouse applications, including human–picker interactions and autonomous shelving transport. While these works underscore the potential of DRL in warehouse environments, their focus and complexity differ significantly from the dynamic order batching and picker routing problem addressed in our study. Collectively, these pioneering efforts (including those discussed earlier in this section) affirm the suitability of DRL for dynamic and complex decision-making in warehouse environments, thereby supporting our methodological choice and underscoring the distinct contribution of our work.
2.3 Broader DRL Applications for Dynamic Assignment and Routing
Our proposed model for order picking aligns with broader applications of DRL in dynamic task allocation and routing, where tasks arrive unpredictably and must be assigned in real time. We highlight here several related studies that, while addressing distinct problem domains, share methodological similarities with our work. For example, da Costa et al. (2023) investigate the traveling maintainer problem, in which a single maintainer responds to dynamically emerging maintenance requests. As with our order picker handling incoming orders, this involves real-time routing, task prioritization, and decision-making under uncertainty. Both studies use DRL to optimize long-term efficiency; however, their objective is to minimize maintenance costs for dispersed assets, whereas we focus on reducing order wait times within a structured warehouse layout.
DRL has also been applied to various other transportation and logistics challenges. Ying et al. (2020) use an actor-critic DRL approach for metro train scheduling under stochastic demand, demonstrating DRL’s effectiveness in large-scale dynamic control. Akkerman et al. (2025) compare RL-based value function approximation and policy function approximation for dynamic vehicle routing with stochastic customer requests, adjusting routes in response to new demand. Similarly, Kitchat et al. (2024) develop a Deep Deterministic Policy Gradient algorithm for distributing epidemic prevention materials during public health emergencies, illustrating the adaptability of DRL to real-time allocation and routing decisions in supply chain logistics. These studies underscore DRL’s versatility in transportation logistics, particularly in vehicle routing and scheduling. In contrast, our approach applies DRL to warehouse order picking, where efficiency gains emerge from optimizing picker movement and batching strategies.
In inventory management, Dehaybe et al. (2024) apply DRL to lot-sizing under non-stationary demand, demonstrating the method’s ability to learn near-optimal replenishment policies without frequent re-optimization. Similarly, Oroojlooyjadid et al. (2022) employ a Deep Q-Network in the Beer Game, showing how DRL can optimize decentralized, multi-stage ordering decisions even under irrational agent behavior. By outperforming base-stock policies and adapting rapidly to changing conditions, their approach illustrates DRL’s flexibility in complex inventory scenarios. Together, these examples underscore DRL’s versatility across diverse supply chain and logistics challenges.
Despite these advancements, existing studies do not provide a comprehensive DRL framework tailored specifically to the warehouse order picking problem. Our contribution lies in modeling warehouse-specific constraints, real-time batching, and picker routing strategies. Further, our reward function explicitly captures key efficiency metrics in warehouse operations, ensuring practical applicability.
@@ 2.4 Research Gap
Table 1: Comparison of Related Order Picking Literature
Reference Order Online Single-Block Minimization Methodology Comments
Batching Routing Warehouse Objective
Layout
Ratliff and Rosenthal (1983) travel distance Exact Solves the static picker routing problem.
Lu et al. (2016) travel distance Exact Orders are batched on a first-come, first-served basis.
Scholz et al. (2016) travel distance Exact+ Heuristic Uses Iterated Local Search for order batching, combined with an exact algorithm for routing optimization.
Aerts et al. (2021) travel distance Heuristic Models the problem as a VRP variant and applies Variable Neighborhood Search to find solutions.
Cals et al. (2021) tardy orders DRL Order sequencing for routing is determined using a heuristic approach.
Beeks et al. (2022) tardy orders & picking cost DRL Extends Cals et al. (2021) to incorporate two objectives.
Dunn et al. (2024) travel distance DRL Solves the static picker routing problem using DRL.
Begnardi et al. (2024) order wait time DRL Grid-based warehouse layout; Focuses on order assignment to multiple pickers to reduce delays; Baselines lack TSP routing or traditional order picking heuristics.
This Study order wait time & travel distance DRL
Our work is the first comprehensive investigation into dynamic order picker routing using DRL, allowing the agent to implicitly learn order batching strategies within a unified learning environment. Table 1 provides a comparative analysis of related order picking literature, highlighting the unique contributions of our study. By integrating assignment and routing, our DRL agent provides rapid, on-the-fly decision-making, crucial for reducing delays in dynamic settings. We prioritize the development of a simple and effective model for single-block warehouses, showing its efficacy in minimizing order processing times, thereby addressing a clear gap in existing research.
@@ 3 Problem Description
Consider the order-picking problem for a single picker within a rectangular warehouse. The warehouse consists of a single block with multiple parallel aisles, interconnected by two cross-aisles located at the ends. Inventory storage locations are positioned on both sides of each aisle. Figure 1 provides a visual representation of this warehouse layout. The picker must retrieve requested items from designated storage locations within the warehouse. Movement is restricted to the aisles, where the picker can access storage locations on either side. However, there is no lateral movement within an aisle. To transition between aisles, the picker must use one of the two cross-aisles. These cross-aisles do not contain any storage locations and serve solely as pathways for aisle transitions. Notably, the picker can only enter or exit an aisle through one of these cross-aisles.
@@
The order-picking device or vehicle utilized for transporting these items has a total capacity of $K$ items. Initially, the picker starts from the depot and moves through the warehouse to collect the assigned items and delivers them to the depot, subject to the capacity constraint at all times. Figure 1 illustrates that this warehouse layout offers two alternative paths the picker can take to collect the next assigned order, denoted by the distances $\rho$. We assume that multiple orders may be received from across the warehouse at any time, with each order representing a single unit and having the same level of priority. Additionally, we assume the requested items are identical in size, a common assumption in the literature (Aerts et al., 2021; Lu et al., 2016; Pansart et al., 2018; Scholz et al., 2016). This assumption is not overly restrictive, as the smallest item in the warehouse can serve as a reference unit (e.g., 1 unit), with larger items represented as multiples of this unit to account for size differences. To prevent fractional item pickups in such scenarios, the proposed method (presented in the next section) can handle this seamlessly without requiring re-training. Specifically, orders can be temporarily excluded from the system if the picker’s capacity is insufficient to pick up the entire order. These excluded orders are reintegrated into the system once the picker completes a drop-off at the depot.
We will discuss the notation details of distances in the next section. The primary objective is to minimize the average waiting time for orders, and thereby increasing the throughput in warehouse operations. The picker route is determined based on the demand forecast, and a dynamically updated pick-list, which is adjusted based on incoming orders and the remaining orders to be collected. This optimization problem is considered over a time horizon that resembles a human picker’s work shift.
A key distinction of our problem setting, compared to the majority of the literature, is that it addresses a dynamic order-picking problem rather than a static one, where predefined pick lists are optimized. In our setting, orders arrive dynamically into the system, and demand fluctuates throughout the work shift. This dynamic environment necessitates real-time decision-making for order batching and traversal routes to minimize order wait times. Reducing these wait times not only enhances customer satisfaction but also frees up resources to fulfill additional orders—an especially critical factor in the rapidly expanding e-commerce sector. Our approach further enhances order picking by incorporating implicit demand forecasting, enabling pickers to anticipate future orders and optimize their paths proactively. This means that instead of taking explicit forecasting steps, our DRL agent implicitly integrates forecasting into its actions. Specifically, the DRL agent’s decisions reflect an understanding of potential future orders, even if they do not align directly with current unfulfilled orders. For example, the agent may opt to wait instead of taking immediate action, despite pending orders in the system. Alternatively, it may navigate to aisles or locations that are currently empty, anticipating that future orders will require those areas. This behavior contrasts sharply with traditional methods, which rely solely on observed order arrivals and re-optimize the system each time a new order appears, without accounting for future demand. The effectiveness of this indirect approach to demand forecasting is demonstrated in our numerical results, which show that the proposed method significantly reduces order wait times compared to traditional methods. These findings underscore the advantages of integrating real-time adaptability and implicit forecasting into dynamic order-picking systems.